HETAV PATEL

Grade 11,

Grade:10

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
We know that:
\frac{sin A}{a}=\frac{sin B}{b}=\frac{sin C}{c} = k(say)
The given equation can be given by:
\frac{a^{2}-b^{2}}{a^{2}+b^{2}}=\frac{sin^{2}A-sin^{2}B}{sin^{2}A+sin^{2}B}
Also,
{sin^{2}A-sin^{2}B}=sin\left ( A+B \right )sin\left ( A-B \right )=sin\left ( C \right )sin\left ( A-B \right )
This proves the required thing.
Regards
Sumit